OpenStudy (anonymous):

given y>0 and dy/dx=3x^2+4x/y. if the point (1,sqrt{10}) is on the graph relating x and y, then what is y when x=0?

myininaya (myininaya):

do you know to find y given y'?

OpenStudy (anonymous):

dy/dx

OpenStudy (anonymous):

?

myininaya (myininaya):

you have to integrate y' to get y

myininaya (myininaya):

oh wait

myininaya (myininaya):

i didn't see that other y

myininaya (myininaya):

is it \[y'=\frac{3x^2+4x}{y} ?\]

OpenStudy (anonymous):

yes

myininaya (myininaya):

ok do you know separation of variables?

OpenStudy (anonymous):

quotient rule product rule ?

myininaya (myininaya):

\[\frac{dy}{dx}=\frac{3x^2+4x}{y}\] you want y so we need to integrate but before we do that we need to separate our y's from our x's

myininaya (myininaya):

\[y dy=(3x^2+4x) dx\]

OpenStudy (anonymous):

oh yes i know how to do that

myininaya (myininaya):

i multiplied y on both sides and I multiplied dx on both sides now we need to integrate both sides

myininaya (myininaya):

\[\int\limits_{}^{}y dy=\int\limits_{}^{}(3x^2+4x) dx\]

OpenStudy (anonymous):

whats that long thing u just put

myininaya (myininaya):

\[\frac{y^2}{2}=\frac{3x^3}{3}+\frac{4x^2}{2}+C\] whenever you integrate you need to do +C

OpenStudy (anonymous):

i havent learned that thing yet

myininaya (myininaya):

its the only way

myininaya (myininaya):

what about finding the antiderivative?

myininaya (myininaya):

do you know what that means?

OpenStudy (anonymous):

i learned that friday

OpenStudy (anonymous):

yes

myininaya (myininaya):

thats with weird symbol means

myininaya (myininaya):

what*

OpenStudy (anonymous):

but i dont think i can use that symbol yet

myininaya (myininaya):

hmmm... that's weird to me

OpenStudy (anonymous):

oh :/

OpenStudy (anonymous):

its fine i guess i should skip that problem and just ask the teacher 2morrow. ill put up a new one.

myininaya (myininaya):

ok but that is all i did was found the antiderivative of both sides with respect to whatever varaible was in d( )

OpenStudy (anonymous):

oh okay thanks either way :)

myininaya (myininaya):

np